OpenStudy (anonymous):

Application of Quadratic Equations Business The manager of a bicycle shop has found that, at a price (in dollars) of p(x)=150-x/5 per bicycle, x bicycles will be sold. a)Find an expression for the total revenue from the sale of x bicycles. (Hint: Revenue = Demand X Price.) b)Find the number of bicycle sales that leads to maximum revenue

OpenStudy (whpalmer4):

Price function is \[P(x) = 150-x/5\] or \[P(x) = \frac{150-x}{5}\]?

OpenStudy (anonymous):

how can i give you a medal

OpenStudy (anonymous):

oh ok.so should i post the second question and tag u to it

OpenStudy (whpalmer4):

this is the question, is it not?

OpenStudy (anonymous):

yes it is

OpenStudy (whpalmer4):

okay. so, I asked for clarification about the form of the price function. which one is it?

OpenStudy (whpalmer4):

I'm guessing perhaps the first one, because the second could be written as 30-x/5, but I've found that people tend not to heed the operator precedence when copying problems to the forum, and it could really be either one.

OpenStudy (anonymous):

dont reallu=y get you clearly.cos i got ma a) as x(150-x/5)=150x -x^2/5

OpenStudy (whpalmer4):

Revenue = number bikes sold (x) * price of bike (150-x/5) = 150x - x^2/5 now we need to find the number of bikes sold that maximizes revenue. Again, this is a parabola...

OpenStudy (anonymous):

yeah i thought as much

OpenStudy (anonymous):

ok with vertex h=-b/2a

OpenStudy (anonymous):

h=x^2/5 /2(150) and this is where i got confused

OpenStudy (whpalmer4):

remember, a and b are the coefficients, and don't include the variable itself.

OpenStudy (whpalmer4):

\[y = ax^2+bx+c\]vertex at \((-b/2a, y(-b/2a))\) \[y = -\frac{1}{5}x^2+150x\]\[a=-\frac{1}{5}, b = 150\] \[-\frac{150}{2*(-\frac{1}{5})} = 150 * \frac{5}{2} = \]

OpenStudy (anonymous):

oh ok. so if im to find the maximum revenue i would need to put 375 in the very first revenue equation we got right? thus150(375)-(375)^2/5

OpenStudy (whpalmer4):

Yes, that should do the trick.

OpenStudy (anonymous):

perfect.